# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ Abstract.
In this paper, we prove several new results on certain ${\mathbb{F}}_{p}$-vector spaces of logarithmic differential forms in characteristic $p$ called spaces $L_{m+1,n}$. Expanding the previous work by the first two authors, we prove positive and negative results for the existence of spaces $L_{m+1,n}$ in many situations, as well as a classification of all spaces $L_{12,2}$ and $L_{15,2}$ for $p=3$. The novel tools used are Moore determinants and computational algebra.
@@ 1. Introduction
Algebraic geometry in positive characteristic is a rich subject that has been intensively studied ever since the foundations of algebraic geometry were rigorously established. [^1] Most notably, studying varieties over finite fields was the main motivation that prompted Weil to write his seminal book Foundations of Algebraic Geometry [1]. As a result, a plethora of positive characteristic methods have been developed. These mathematical tools not only unlock novel insights into the geometry of algebraic varieties but also have surprising applications to other disciplines. For example, they form the foundations for the development of advanced coding techniques and cryptographic protocols used nowadays. Moreover, the richness of algebraic geometry in positive characteristic has been also exploited to prove results in characteristic zero. Roughly speaking, this is made possible by the use of two complementary procedures: lifting to characteristic 00, that assigns an object in characteristic zero with a given object in positive characteristic, and reduction modulo $p$, that assigns an object in positive characteristic with a given object in characteristic zero.
@@
Given $k$ an algebraically closed field of characteristic $p>0$ and strictly positive integers $n,m\in{\mathbb{N}}$, a set $\Omega$ of differential forms on ${\mathbb{P}}^{1}_{k}=k\cup\infty$ is called a space $L_{m+1,n}$ if it satisfies the following conditions
@@
  • Every $\omega\in\Omega-\{0\}$ has a unique zero at $\infty$ of order $m-1$.
Classically, the motivation for studying spaces $L_{m+1,n}$ arose from the following lifting problem: let $G\subset\operatorname{Aut}_{k}k[\![z]\!]$ be a finite order subgroup of $k$-automorphisms of the ring $k[\![z]\!]$ of formal power series over $k$. It is said that $G$ lifts to characteristic zero if there exists a finite extension $R$ of the ring of Witt vectors $W(k)$ with uniformizer $\pi$ and a commutative diagram
(1.1) ${\mathrm{Aut}_{R}R[\![Z]\!]}$${G}$${\mathrm{Aut}_{k}k[\![z]\!]}$$\scriptstyle{r}$$\scriptstyle{j}$$\scriptstyle{i}$
where $i$ is the inclusion and $r$ is the reduction modulo $\pi$. The local lifting problem for curves is the question of determining which embeddings $G\subset\operatorname{Aut}_{k}k[\![z]\!]$ admit a lifting. This problem has now a long tradition and can be studied from many angles: the interested reader is referred to the surveys [2] and [3] for a thorough discussion. In this context, the spaces $L_{m+1,n}$ arise in the elementary abelian case, that is, when $G\cong\leavevmode\nobreak\ ({\mathbb{Z}}/p{\mathbb{Z}})^{n}$. This connection is now well established, thanks to the results obtained in [4], [5] (for $n=1$), [6] and [7] (for $n>1$). Thanks to these investigations, we know that the existence of a space $L_{m+1,n}$ for a given triple $(p,n,m)$ is equivalent to the existence of a lifting of $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$ with equidistant branch locus and $m+1$ fixed points. [^2] See [8, Théorème 11] for the construction of a lifting with equidistant branch locus given a space $L_{m+1,n}$. We also know that $m+1$ is necessarily of the form $\lambda p^{n-1}$ for some $\lambda>0$, a fact that leads naturally to the following question
@@
In this paper we make progress on this largely open question, as well as introduce tools that can be used to see it in a new light. Let us recall what was known prior to this paper: if $n=2$ and $p=2$ there are spaces $L_{2\lambda,2}$ for every $\lambda>0$ ([7, Théorème 2.2.4]). If $n=2$ and $p\geq 3$ then there are no spaces $L_{p,2}$ and no spaces $L_{3p,2}$, while spaces $L_{2p,2}$ exist if, and only if $p=3$ ([7, Théorème 2.2.5]). If $n=2$ and $\lambda\geq 4$ the only examples of spaces $L_{\lambda p,2}$ that were known before this paper have the special property that $\lambda$ is a multiple of $p-1$ and obey a strict geometric constraint. Similarly, in the case $n\geq 3$ the known examples (see Section 5.1) satisfy $p-1|\lambda$ and are of a very special nature.
Our contributions to the question above solve the problem of existence of spaces $L_{\lambda p^{n-1},n}$ in three distinct cases:
@@
More precisely, we provide a complete classification of spaces $L_{\lambda p,2}$ for $p=3$ and $\lambda=4,5$, we show that for $p>3\lambda$ there are no spaces $L_{\lambda p,2}$. Moreover, we prove new results on the existence of spaces $L_{\lambda p^{n-1},n}$ for $n\geq 3$ that, when applied to the case $p=2$, lead to construction of large classes of spaces $L_{\lambda 2^{n-1},n}$ when $\lambda$ is either even or congruent to $1$ modulo $22^{n}-$ 2. In proving the results above, we make use of techniques that have not been exploited in this context before, and that we believe to be of independent interest, such as computational commutative algebra, étale pullbacks, and Moore determinants. The first main result presented in the paper is the case $n=2$ and $p>3\lambda$.
@@
The proof of this theorem relies on a result of Pagot [^3] Proposition 3.1, first appearing as [8, Proposition 7], that shows how the existence of $\Omega$ a space $L_{\lambda p,2}$ for fixed $\lambda$ and $p$ is equivalent to the existence of two polynomials $Q_{1},Q_{2}\in k[X]$ of degree $\lambda$ satisfying three conditions. The first condition is that $Q_{1}$ and $Q_{2}$ are ${\mathbb{F}}_{p}$-linearly independent. The second condition is that the set of poles of $\Omega$ coincides with the set of zeroes of the polynomial $Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}$. The third condition can be expressed as a multivariate polynomial system, whose indeterminates are the coefficients of $Q_{1}$ and $Q_{2}$. This system is in general very complicated, but a subsystem of necessary conditions can be extracted thanks to the fact that the nonzero elements of $\Omega$ are logarithmic and hence their poles have residues in ${\mathbb{F}}_{p}^{\times}$. In the context of our theorem, the necessary conditions can be expressed in terms of the coefficients of $Q_{1}$ and $Q_{2}$ and become tractable enough to get a contradiction under the assumptions of the theorem.
@@
In the case $p>3$ the above simplifications are not enough to conclude. However, a finer strategy can be employed to obtain a classification of all spaces $L_{4p,2}$ for all prime $p$’s, even though it requires substantial more work. In fact, in this case Theorem 3.7 allows us to consider only the cases $p=3$ (discussed above) and $p=5,7,11$. In these cases, we can use a combination of elementary arguments to find previously undiscovered relations between the zeroes of $Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}$. In almost all cases (i.e. except when $Q_{1}$ and $Q_{2}$ are of a very specific form) these relations take the form of vanishing of certain Schur polynomials, that can be turned, using Jacobi-Trudi relations, into the vanishing of certain Toeplitz determinants involving symmetric polynomials in some of the zeroes. These new relations are easier to work out and show that, for $p=5$ all the spaces $L_{20,2}$ are of a previously known form, and for $p=7,11$ there are no spaces $L_{4p,2}$. A discussion of the remaining outstanding case necessitates additional arguments and the computation of Grobner bases, but can be achieved and confirms that there are no spaces $L_{4p,2}$ other than the previously known cases in general. The write up of this result is in preparation.
We then turn our attention to the case of spaces $L_{\lambda p^{n-1},n}$. The crucial new tool that we use in this setting is the Moore determinant $\Delta_{n}(\underline{a})$ of a $n$-uple of elements $\underline{a}:=(a_{1},\dots,a_{n})$ in a field of characteristic $p$. This is defined as the determinant of the associated Moore matrix, namely we have
@@
The Moore determinant is a basic object in arithmetic in characteristic $p$, due to its relationship with the theory of additive polynomials (cf. [9, Chapter 1]). In this paper, we apply a mix of classical and recent results (obtained in [10]) on Moore determinants to describe effectively the differential forms belonging to a space $L_{\lambda p^{n-1},n}$. Our exposition of Moore determinants is self-interested and self-contained: we put all and only the results we need in the dedicated Appendix A, and include there further, more complete, references for the interested reader.
Moore determinants allow us to prove two main new results about spaces $L_{\lambda p^{n-1},n}$. These take the form of a necessary and sufficient criterion for the existence of spaces $L_{\lambda p^{n-1},n}$, which for $n=2$ boils down to Proposition 3.1.
Theorem (cf. Theorems 4.7 and 4.8).
Let $n\geq 2$ and let $\Omega$ be a $n$-dimensional ${\mathbb{F}}_{p}$-vector space of differential forms on ${\mathbb{P}}^{1}_{k}$, generated by elements $\omega_{1},\dots,\omega_{n}$. Then $\Omega$ is a space $L_{\lambda p^{n-1},n}$ if, and only if, the following conditions are met:
  1. (i)
    There exists a non-zero logarithmic form $\omega\in\Omega$.
  2. (ii)
    There exist a $n$-uple of polynomials $\underline{Q}:=(Q_{1},\dots,Q_{n})\in k[X]^{n}$ such that
    • The polynomials $Q_{i}$ are of degree $\lambda$ for every $i=1,\dots,n$.
    • The leading coefficients $q_{1},\dots,q_{n}$ of $Q_{1},\dots,Q_{n}$ satisfy $\Delta_{n}(q_{1},\dots,q_{n})\neq 0$.
@@
      for every $i=1,\dots,n$, where $\underline{\widehat{Q_{i}}}$ denotes the $n-1$-uple obtained by removing $Q_{i}$ from $\underline{Q}$.
This condition simplifies our task of classifying spaces $L_{\lambda p^{n-1},n}$ insofar as it gives us a practical recipe to use the polynomials $Q_{i}$’s to build a basis of $\Omega$. In Section 4.2 we develop a constructive strategy to build suitable $Q_{i}$’s, that is fruitfully applied in the case $p=2$ to create a large class of new examples. Our main result in this sense is the following:
@@
Let $\lambda$ be either even or $\lambda\equiv 1\mod(2^{n}-2)$. Then there exist infinitely many equivalence classes of spaces $L_{\lambda 2^{n-1},n}$.
Then, we turn our attention to the examples of spaces $L_{\lambda p^{n-1},n}$ for $n\geq 3$ that were known in the literature prior to the present paper. We remark that all these spaces share a very special structure: each of them is an étale pullback of a space $L_{(p-1)p^{n-1},n}$ whose set of poles is the set of nonzero elements in a $n$-dimensional ${\mathbb{F}}_{p}$-subvector space of $k$. We call standard any space $L_{(p-1)p^{n-1},n}$ whose poles satisfy the property above, and we apply Theorems 4.7 and 4.8 in the special case of standard spaces, shedding new light on their arithmetic properties. As a result, we are able to characterize the subspaces of standard spaces as étale pullbacks of standard spaces of lower dimension.
@@
Let $\Omega$ be a standard $L_{(p-1)p^{n-1},n}$ space and let $\widetilde{\Omega}$ be a proper $r$-dimensional subspace of $\Omega$. Then, there exists a standard $L_{(p-1)p^{r-1},r}$ space $\Omega^{\prime}$ and a degree $p^{n-r}$ morphism $\sigma:{\mathbb{P}}^{1}_{r}\to{\mathbb{P}}^{1}_{r}$ ramified only at $\infty$ such that $\widetilde{\Omega}$ is the pullback of $\Omega^{\prime}$ under $\sigma$.
@@
In Section 2, we present known results and useful constructions on $L_{\lambda p^{n-1},n}$ that are used in all the other sections. In Section 3, we recall an important characterization of spaces $L_{\lambda p,2}$, due to Pagot (Proposition 3.1), and we prove the non existence of spaces $L_{\lambda p,2}$ when $p>3\lambda$. In Section 4, we generalize Proposition 3.1 to the case of spaces $L_{\lambda p^{n-1},n}$ and we apply this to the case $p=2$ to construct our new examples of spaces $L_{\lambda 2^{n-1},n}$. In Section 5, we define and study standard $L_{\lambda p^{n-1},n}$-spaces. In Section 6, we fix $p=3$ and provide a complete classification of spaces $L_{12,2}$ and $L_{15,2}$. Finally, in Appendix A we collect all the results on Moore determinants that are used throughout the paper (mostly in Sections 4 and 5).
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. Recall that a differential form on ${\mathbb{P}}^{1}_{k}$ can be written as $\omega=f(X)dX$ for $f(X)\in k(X)$, the field of rational fractions in one variable over $k$. Such a differential form is called logarithmic if it is of the form $\omega=\frac{dF}{F}$ for some $F\in k(X)$. We usually denote by $\Omega$ a space $L_{m+1,n}$ and by $\{\omega_{1},\dots,\omega_{n}\}$ a basis for this space. For a given subset $S\subset\Omega$, we denote by ${\mathcal{P}}(S)$ the subset of $k$ consisting of elements that are poles of at least a non-zero differential form in $S$. If $\omega\in\Omega-\{0\}$, we write ${\mathcal{P}}(\omega)$ for the set of poles of $\omega$. One deduces from the definition that the set ${\mathcal{P}}(\omega)$ consists of $m+1$ simple poles. Finally, given a finite set of logarithmic differential forms $\{\omega_{1},\dots,\omega_{n}\}$, we denote by $\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ the ${\mathbb{F}_{p}}$-vector space that they generate. Most of the times, this will not be a space $L_{m+1,n}$, but it will be one under certain predetermined conditions.

### v2

@@ Abstract.
Let $p$ be a prime number. Motivated by the local lifting problem for $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$ with $n>1$, we prove several new results on certain ${\mathbb{F}}_{p}$-vector spaces of logarithmic differential forms on the projective line in characteristic $p$, called spaces $L_{m+1,n}$. Expanding the previous work by the first two authors, we prove positive and negative results for the existence of spaces $L_{m+1,n}$ in many situations. Moreover, we classify all spaces $L_{4p,2}$ for any $p$, and all spaces $L_{15,2}$ for $p=3$. Among the novel tools we use, Moore determinants and computational algebra play a prominent role.
@@ 1. Introduction
Algebraic geometry in positive characteristic is a rich subject that has been intensively studied ever since the foundations of algebraic geometry were rigorously established. [^1] Most notably, studying varieties over finite fields was the main motivation that prompted Weil to write his seminal book Foundations of Algebraic Geometry [Weil46]. Weil’s theory was subsequently generalized extensively by Grothendieck into the theory of schemes, which forms the foundation of modern algebraic geometry. As a result, a plethora of positive characteristic methods have been developed. These mathematical tools not only unlock novel insights into the geometry of algebraic varieties but also have surprising applications to other disciplines. For example, they form the foundations for the development of advanced coding techniques and cryptographic protocols used nowadays. Moreover, the richness of algebraic geometry in positive characteristic has been also exploited to prove results in characteristic zero. Roughly speaking, this is made possible by the use of two complementary procedures: lifting to characteristic 0, that assigns an object in characteristic zero with a given object in positive characteristic, and reduction modulo $p$, that assigns an object in positive characteristic with a given object in characteristic zero.
@@
Given $k$ an algebraically closed field of characteristic $p>0$ and strictly positive integers $n,m\in{\mathbb{N}}$, a set $\Omega$ of differential forms on the curve ${\mathbb{P}}^{1}_{k}$ is called a space $L_{m+1,n}$ if it satisfies the following conditions
@@
  • Every $\omega\in\Omega-\{0\}$ has a unique zero, which is at $\infty$ and has order m1m-1.
Classically, the motivation for studying spaces $L_{m+1,n}$ arose from the following local lifting problem:
Let $G$ be a finite group, and k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!] a $G$-galois extension. Does there exist a finite extension $R$ of the ring of Witt vectors $W(k)$ with uniformizer $\pi$ and a $G$-galois extension R[[Z]]/R[[T]]R[\![Z]\!]/R[\![T]\!] such the $G$ action on R[[Z]]R[\![Z]\!] reduces modulo $\pi$ to the given $G$ action on k[[z]]k[\![z]\!]?
If this problem has a solution for a $G$-extension k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!], we say that the extension lifts to characteristic 0 and that R[[Z]]/R[[T]]R[\![Z]\!]/R[\![T]\!] is a lift of the extension. In this case, we can consider the cover $f:\penalty 10000\ \mathrm{Spec}(R[\![Z]\!])\longrightarrow\mathrm{Spec}(R[\![T]\!])$: it is unramified at the prime ideal $(\pi)$ and the branch points of R[[Z]]/R[[T]]R[\![Z]\!]/R[\![T]\!] are defined as the divisors $b$ of $\mathrm{Spec}(R[\![T]\!])$ such that $f$ is ramified at $f^{-1}(b)$. After extending $R$, we can assume that all branch points of this cover are $R$-rational. Then, the set of these branch points is also called the branch locus of R[[Z]]/R[[T]]R[\![Z]\!]/R[\![T]\!].
The local lifting problem has a relatively long tradition and can be studied from many angles: the interested reader is referred to the surveys [Obus12a] and [Obus19] for a thorough discussion. In this paper we are interested in the elementary abelian case, that is when $G\cong({\mathbb{Z}}/p{\mathbb{Z}})^{n}$. When $n=2$, [GreenMatignon98, Theorem 5.1] gives necessary and sufficient conditions for lifting, involving the ramification filtration and combinatorics of the branch locus of lifts of intermediate p-cyclic extensions. This was recently generalized by Yang [Yang25] to the case $n>2$, and was used as a guide to build lifts for some $({\mathbb{Z}}/2{\mathbb{Z}})^{3}$-extensions. Subsequent work of Pagot [Pagot25], combining Yang’s results and the notion of Hurwitz trees, goes further, solving the local lifting problem for $p=2$ and $n=2,3$.
The investigation of spaces $L_{m+1,n}$ of the present paper results in necessary and sufficient conditions for the existence of lifts of $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$-extensions with branch locus of a very specific type, for arbitrary $p$ and $n>1$. In fact, thanks to the results obtained in [GreenMatignon99], [HenrioThese] (for $n=1$), [Matignon99] and [PagotThese] (for $n>1$), we know that the existence of a space $L_{m+1,n}$ for a given triple $(p,n,m)$ is equivalent to the existence of a lift of a $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$– extension k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!] with a unique ramification jump of conductor $m$ and an equidistant branch locus, that is, the branch locus $\{(T-t_{i})\in\mathrm{Spec}(R[\![T]\!]):i\in I\}$ is such that the set $\{v_{R}(t_{i}-t_{j}):i,j\in I,i\neq j\}$ consists of a unique element. Up to change of the parameter $T$, we can always assume that there exists a unique integer $r>0$ such that $v_{R}(t_{i}-t_{j})=v_{R}(t_{i})=r$ for every $i,j\in I$ with $i\neq j$. Given such a lift, one can build a space $L_{m+1,n}$ whose set of poles is $\{\frac{t_{i}}{\pi^{r}}\mod\pi:0\leq i\leq m\}\subset k$. Conversely, for every space $L_{m+1,n}$, a construction by Pagot ([Pagot02, Théorème 11]) shows the existence of a lift with a unique ramification jump of conductor $m$ and an equidistant branch locus that induces the given space $L_{m+1,n}$. Moreover, for such a lift to exist, we know that $m+1$ is necessarily of the form $\lambda p^{n-1}$ for some $\lambda>0$, a fact that leads naturally to the following question
@@
In this paper we make progress on this largely open question, as well as introduce tools that can be used to see it in a new light. Let us recall what was known prior to this paper: if $n=2$ and $p=2$ there are spaces $L_{2\lambda,2}$ for every $\lambda>0$ ([Pagot02, Théorème 8]). If $n=2$ and $p\geq 3$ then there are no spaces $L_{p,2}$ and no spaces $L_{3p,2}$, while spaces $L_{2p,2}$ exist if, and only if $p=3$ ([Pagot02, Théorème 9]). If $n=2$ and $\lambda\geq 4$ the only examples of spaces $L_{\lambda p,2}$ that were known before this paper have the special property that $\lambda$ is a multiple of p1p-1 and obey a strict geometric constraint. Similarly, in the case $n\geq 3$ the known examples (see Section 5.1) satisfy $p-1|\lambda$ and are of a very special nature.
Our contributions to the question above solve the problem of existence of spaces $L_{\lambda p^{n-1},n}$ in four cases:
@@
More precisely, we provide a complete classification of spaces $L_{\lambda p,2}$ for $p=3$ and $\lambda=4,5$, we show that for $p>3\lambda$ there are no spaces $L_{\lambda p,2}$, and we show that spaces $L_{4p,2}$ exist only for $p\leq 5$, and that for $p=5$ they all are of a very special kind. Moreover, we prove new results on the existence of spaces $L_{\lambda p^{n-1},n}$ for $n\geq 3$ that, when applied to the case $p=2$, lead to construction of large classes of spaces $L_{\lambda 2^{n-1},n}$ when $\lambda$ is either even or congruent to $1$ modulo $2^{n}-2$. In proving the results above, we make use of techniques that have not been exploited in this context before, and that we believe to be of independent interest, such as computational commutative algebra, étale pullbacks, and Moore determinants. The first main result presented in the paper is the case $n=2$ and $p>3\lambda$.
@@
The proof of this theorem relies on a result of Pagot [^2] Proposition 3.1, first appearing as [Pagot02, Proposition 7], that shows how the existence of $\Omega$ a space $L_{\lambda p,2}$ for fixed $\lambda$ and $p$ is equivalent to the existence of two polynomials $Q_{1},Q_{2}\in k[X]$ of degree $\lambda$ satisfying three conditions. The first condition is that $Q_{1}$ and $Q_{2}$ are ${\mathbb{F}}_{p}$-linearly independent. The second condition is that the set of poles of $\Omega$ coincides with the set of zeroes of the polynomial $Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}$. The third condition can be expressed as a multivariate polynomial system, whose indeterminates are the coefficients of $Q_{1}$ and $Q_{2}$. This system is in general very complicated, but a subsystem of necessary conditions can be extracted thanks to the fact that the nonzero elements of $\Omega$ are logarithmic and hence their poles have residues in ${\mathbb{F}}_{p}^{\times}$. In the context of our theorem, the necessary conditions can be expressed in terms of the coefficients of $Q_{1}$ and $Q_{2}$ and become tractable enough to get a contradiction under the assumptions of the theorem.
@@
In the case $p>3$ the above simplifications are not enough to conclude. However, a finer strategy can be employed to obtain a classification of all spaces $L_{4p,2}$ for all prime $p$’s, even though it requires more work. In this case Theorem 3.9 allows us to consider only the cases $p=3$ (discussed above) and $p=5,7,11$, which can be solved through a combination of elementary arguments and the support of computational methods, finally yielding the following result.
Theorem (cf. Theorem 7.1).
Let $p$ be a prime number and let $\Omega$ be a space $L_{4p,2}$. Then $p\in\{2,3,5\}$ and if $p=5$ then we obtain a full classification of spaces $L_{20,2}$.
The proof of this result consists in splitting the analysis of the coefficients of the polynomials $Q_{1},Q_{2}$ into three subcases. In each of these subcases we show that $Q_{1}$ and $Q_{2}$ are necessarily biquadratic polynomials. Then, either by solving the polynomial system given by Proposition 3.1 via Grobner bases computations or, where possible, by explicit computing certain weighted power sums in the zeroes of $Q_{1}+jQ_{2}$ and showing that they must satisfy a certain relation, we find a contradiction in all cases unless $p=5$ and $\Omega$ is of a very special kind (the one considered in Section 5.1).
We then turn our attention to the case of spaces $L_{\lambda p^{n-1},n}$. The crucial new tools that we need in this context are Moore determinants and Dickson invariants. The Moore determinant $\Delta_{n}(\underline{a})$ of a $n$-tuple of elements $\underline{a}:=(a_{1},\dots,a_{n})$ in a field of characteristic $p$ is defined as the determinant of the associated Moore matrix, namely we have
@@
The Moore determinant is a basic object in arithmetic in characteristic $p$, due to its relationship with the theory of additive polynomials (cf. [Goss96, Chapter 1]). In this paper, we apply a mix of classical and recent results (obtained in [FresnelMatignon23]) on Moore determinants to describe effectively the differential forms belonging to a space $L_{\lambda p^{n-1},n}$. Moore determinants are also important to define the Dickson invariants $c_{n,i}\in{\mathbb{F}}_{p}[X_{1},\dots,X_{n}]$. More precisely, for a fixed $n$ and $0\leq i\leq n$, one defines these by the formula
$\frac{\Delta_{n+1}(\underline{X},T)}{\Delta_{n}(\underline{X})}=\sum_{i=0}^{n}(-1)^{n-i}c_{n,i}T^{p^{i}}.$
Dickson invariants play a central role in the invariant theory of finite groups and were defined by Dickson [Dickson11] to prove that the invariant algebra ${\mathbb{F}}_{p}[X_{1},\dots,X_{n}]^{GL_{n}({\mathbb{F}}_{p})}$ is a polynomial algebra generated by the $c_{n,i}$’s. Our exposition of Moore determinants and Dickson invariants is self-interested and self-contained: we introduce all and only the results we need in the dedicated Appendix A and include there further references for the keen reader.
When applied to spaces $L_{\lambda p^{n-1},n}$, these techniques prove to be very fruitful, allowing to prove a new condition for the existence of spaces $L_{\lambda p^{n-1},n}$ which for $n=2$ boils down to Proposition 3.1.
Theorem (combining Theorems 4.7, 4.10 and A.6).
Let $n\geq 2$ and let $\Omega$ be a $n$-dimensional ${\mathbb{F}}_{p}$-vector space of differential forms on ${\mathbb{P}}^{1}_{k}$, generated by elements $\omega_{1},\dots,\omega_{n}$. Then $\Omega$ is a space $L_{\lambda p^{n-1},n}$ if, and only if, there exist a $n$-tuple of polynomials $\underline{Q}:=(Q_{1},\dots,Q_{n})\in k[X]^{n}$ of degree $\lambda$ such that
  • The leading coefficients $q_{1},\dots,q_{n}$ of $Q_{1},\dots,Q_{n}$ satisfy $\Delta_{n}(\underline{q})\neq 0$.
@@
    for every $i=1,\dots,n$, where $\underline{\widehat{Q_{i}}}$ denotes the n1n-1-tuple obtained by removing $Q_{i}$ from $\underline{Q}$.
  • The n1n-1th Dickson invariant evaluated at the $n$-tuple $\underline{Q}$ satisfies the equation
    $c_{n,n-1}(\underline{Q})^{(p-1)}=-1,$
    where the exponent (p1)(p-1) denotes the p1p-1-th derivative.
This theorem allows us to approach the classification of spaces $L_{\lambda p^{n-1},n}$ from a better viewpoint. Namely, we can start with a $n$-tuple of polynomials $\underline{Q}$ and then check if this satisfies the conditions of the theorem. This approach leads to advances in the classification of spaces $L_{\lambda p^{n-1},n}$ thanks to the fact that the condition $c_{n,n-1}(\underline{Q})^{(p-1)}=-1$ is equivalent to a condition involving the products of two determinants (see Proposition 4.20). In this way, we can achieve new results in the case $p=2$ (for every $n$ and $\lambda$) and in the case $\lambda=1$ (for every $n$ and $p\neq 2$). These are the next two theorems.
@@
Let $p=2$ and $\lambda$ be either even or $\lambda\equiv 1\mod(2^{n}-2)$. Then there exist infinitely many equivalence classes of spaces $L_{\lambda 2^{n-1},n}$.
Theorem (Theorem 4.31).
Let $p>2$. Then, there exist no space $L_{p^{n-1},n}$.
Then, we turn our attention to the examples of spaces $L_{\lambda p^{n-1},n}$ for $n\geq 3$ that were known in the literature prior to the present paper. We remark that all these spaces share a very special structure: each of them is an étale pullback of a space $L_{(p-1)p^{n-1},n}$ whose set of poles is the set of nonzero elements in a $n$-dimensional ${\mathbb{F}}_{p}$-subvector space of $k$. We call standard any space $L_{(p-1)p^{n-1},n}$ whose poles satisfy the property above, and we apply Theorems 4.7 and 4.10 in the special case of standard spaces, shedding new light on their arithmetic properties. As a result, we are able to characterize the proper subspaces of standard spaces as étale pullbacks of standard spaces of lower dimension.
@@
Let $\Omega$ be a standard $L_{(p-1)p^{n-1},n}$ space and let $\widetilde{\Omega}$ be a proper $r$-dimensional subspace of $\Omega$. Then, there exists a standard $L_{(p-1)p^{r-1},r}$ space $\Omega^{\prime}$ and a degree $p^{n-r}$ morphism $\sigma:{\mathbb{P}}^{1}_{k}\to{\mathbb{P}}^{1}_{k}$ ramified only at $\infty$ such that $\widetilde{\Omega}$ is the pullback of $\Omega^{\prime}$ under $\sigma$.
@@
In Section 2, we present known results and useful constructions on $L_{\lambda p^{n-1},n}$ that are used in all the other sections. In Section 3, we recall an important characterization of spaces $L_{\lambda p,2}$, due to Pagot (Proposition 3.1), and we prove the non existence of spaces $L_{\lambda p,2}$ when $p>3\lambda$. In Section 4, we generalize Proposition 3.1 to the case of spaces $L_{\lambda p^{n-1},n}$ and we apply this to the case $p=2$ to construct our new examples of spaces $L_{\lambda 2^{n-1},n}$. In Section 5, we define and study standard $L_{\lambda p^{n-1},n}$-spaces. In Section 6, we fix $p=3$ and $n=2$ and provide a complete classification of spaces $L_{12,2}$ and $L_{15,2}$. In Section 7, we fix $\lambda=4$ and $n=2$ and we deal with the outstanding cases in the classification of spaces $L_{4p,2}$. Finally, in Appendix A we collect all the results on Moore determinants that are used throughout the paper (mostly in Sections 4 and 5).
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. A meromorphic differential form on ${\mathbb{P}}^{1}_{k}$ can be written as $\omega=f(X)dX$ for $f(X)\in k(X)$, the field of rational fractions in one variable over $k$. Given a point $a\in{\mathbb{P}}^{1}_{k}(k)=k\cup\{\infty\}$ and a differential form $\omega=f(X)dX$ on ${\mathbb{P}}^{1}_{k}$, the order of $\omega$ at $a$ is equal to the order $\mathrm{ord}_{a}(f(X))$ if $a\neq\infty$, and it is equal to $\mathrm{ord}_{0}(f(\frac{1}{X}))-2$ if $a=\infty$. When $\mathrm{ord}_{a}(\omega)>0$, the point $a$ is a zero of $\omega$, and when $\mathrm{ord}_{a}(\omega)<0$, the point $a$ is a pole of $\omega$.
A differential form on ${\mathbb{P}}^{1}_{k}$ is called logarithmic if it is of the form $\omega=\frac{dF}{F}$ for some $F\in k(X)$. We usually denote by $\Omega$ a space $L_{m+1,n}$ and by $\{\omega_{1},\dots,\omega_{n}\}$ a basis for this space. For a given subset $S\subset\Omega$, we denote by ${\mathcal{P}}(S)$ the subset of $k$ consisting of elements that are poles of at least a non-zero differential form in $S$. If $\omega\in\Omega-\{0\}$, we write ${\mathcal{P}}(\omega)$ for the set of poles of $\omega$. One deduces from the definition that the set ${\mathcal{P}}(\omega)$ consists of $m+1$ simple poles. Finally, given a finite set of logarithmic differential forms $\{\omega_{1},\dots,\omega_{n}\}$, we denote by $\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ the $\mathbb{F}_{p}$}-vector space that they generate. Most of the times, this will not be a space $L_{m+1,n}$, but it will be one under certain predetermined conditions.
Acknowledgments
Thanks to Jean Fresnel’s careful reading and insightful comments, this paper is better than it otherwise would have been. We dedicate it to his memory.